Find an equation of the line containing the given pair points.
(-3,-8) and (-9,-5)
the equation of the line in slope-intercept form is y =____________
(Simplify your answer. Use integers or fractions for any numbers in the expression)

Answer by confusedgirl81(14)

First we have to find the slope of a line by using the two given points.
x1 y1 x2 y2
(-3,-8) and (-9,-5)

m = y2-y1/x2-x1
-5-(-8)/-9 - (-3) {substitute y2,y1,and x2,x1}
-5+8/-9+3
3/-6 {simplify}
-1/2
so m = -1/2
Now we will use the point slope form to find the equation of a line.
You can use either one of the points already given. I will use the first point. (-3,-8).

y-y1=m(x-x1) {point slope form}
y-(-8) = (-1/2)(x-(-3)) {substitute m,x1,and y1}
y + 8 = -1/2x - 3/2 {subtract 8 from both sides}
y = -1/2x - 3/2 -8 {simplify -3/2-8}
y = -1/2x - 9/2
You can check my take the point slope form and using the other point (-9,-5). You will still end up with the same answer.
So y = -1/2x -9/2 is the equation of the line with a slope of -1/2x that passes through the points (-3,-8) and (-9,-5).
If you need more help, just ask!